Find the x and y intercepts
2x-3y=9
Translate to an equation
B $3 bows and R $5 roll of wrapping paper total $45

Answer by mananth(16946)

2x-3y=9
X intercept
plug y=0
2x-0=9
2x=9
/2
x=9/2
Y intercept
plug x=0
-3y=9
/-3
y=-3
--------------
bows be x numbers
and roll be y numbers
3x+5y=45